An ordinary (fair) die is a cube with the numbers 1 through 6 on the sides (represented by painted spots). Imagine that such a die is rolled twice in succession and that the face values of the two rolls are added together. This sum is recorded as the outcome of a single trial of a random experiment. 


Compute the probability of each of the following events. 

 Event A: The sum is greater than 7. 

 Event B: The sum is divisible by 6. 
Answer #1

a)P(A) =P(sum is greater than 7) =15/36=5/12 (since 15 such outcomes are there out of 36)

b)P(B) =P(sum is 6 ,12) =(5+1)/36 =1/6